OpenStudy (rikkie):

g(x) = 2^(x - 1) - 2 Find the x-intercept I know the x-intercept is when y is zero, but I'm still unclear on the steps. Can someone help me out?

OpenStudy (phi):

in this case g(x) is y, and you g(x)=0 \[ 2^{(x-1)} -2 = 0 \] any idea what the next step should be ?

OpenStudy (rikkie):

Eek, I have a feeling that I do but I'm not sure!

OpenStudy (phi):

add 2 to both sides. can you do that ?

OpenStudy (rikkie):

I did and got 2 when I added 0+2, is that it?

OpenStudy (phi):

like this \[ 2^{(x-1)} -2+2 = 0+2\]

OpenStudy (rikkie):

Oooh I see, I added +2 to the first 2 in it!

OpenStudy (phi):

on the left side -2 + 2 simplifies to 0 on the right side 0+2 is 2 \[ 2^{(x-1)} = 2\]

OpenStudy (phi):

2 can be written as \( 2^1 \) so we could write the equation as \[ 2^{(x-1)} = 2^1\]

OpenStudy (phi):

when the bases are equal (and here the base is 2 on both sides), you can set the exponents equal.

OpenStudy (rikkie):

Thank you for clearing that bit up!

OpenStudy (phi):

can you finish ?

OpenStudy (rikkie):

I think I need a little help finishing, but I definitely understand the steps now

OpenStudy (phi):

set the exponents equal x-1 = 1 What is a good thing to add (to both sides) that makes the -1 next to the x add up to 0?

OpenStudy (rikkie):

1?

geerky42 (geerky42):

Yes.

OpenStudy (rikkie):

Thank you both!

OpenStudy (phi):

you should get x-1+1 = 1+1 or after simplifying x = 2